# Alleviating Hallucinations in Large Vision-Language Models through Hallucination-Induced Optimization

### v1

@@ Abstract
Although Large Visual Language Models (LVLMs) have demonstrated exceptional abilities in understanding multimodal data, they invariably suffer from hallucinations, leading to a disconnect between the generated text and the corresponding images. Almost all current visual contrastive decoding methods attempt to mitigate these hallucinations by introducing visual uncertainty information that appropriately widens the contrastive logits gap between hallucinatory and targeted ones. However, due to uncontrollable nature of the global visual uncertainty, they struggle to precisely induce the hallucinatory tokens, which severely limits their effectiveness in mitigating hallucinations and may even lead to the generation of undesired hallucinations. To tackle this issue, we conducted the theoretical analysis to promote the effectiveness of contrast decoding. Building on this insight, we introduce a novel optimization strategy named Hallucination-Induced Optimization (HIO). This strategy seeks to amplify the contrast between hallucinatory and targeted tokens relying on a fine-tuned theoretical preference model (i.e., Contrary Bradley-Terry Model), thereby facilitating efficient contrast decoding to alleviate hallucinations in LVLMs. Extensive experimental research demonstrates that our HIO strategy can effectively reduce hallucinations in LVLMs, outperforming state-of-the-art methods across various benchmarks.
@@ 1 Introduction
The recent success of Large Vision-Language Models (LVLMs) marks a major milestone in artificial intelligence research OpenAI (2023); Alayrac et al. (2022); Li et al. (2023a); Liu et al. (2023d); Zhu et al. (2023); Bai et al. (2023); Dai et al. (2023); Wang et al. (2023b); Driess et al. (2023). By seamlessly integrating visual cues with Large Language Models (LLMs), LVLMs have demonstrated unparalleled expertise in multimodal comprehension, logical reasoning, and interactive engagement. This integration has ushered in a new era in AI, breaking through traditional limitations and enabling a more holistic understanding of complex information OpenAI (2023); Yang et al. (2023); Lu et al. (2023). Despite these advancements, certain challenges remain, particularly the issue of hallucination Li et al. (2023b); Gunjal et al. (2023); Liu et al. (2023b); Lovenia et al. (2023). Hallucination occurs when the language model generates content that deviates from the image’s actual content, including imagined objects, fabricated scenes, incorrect spatial relationships, and misidentified categories.
Figure 1: (Left) Challenges and Solutions of Contrast Decoding Strategy. Visual Contrastive Decoding, despite introducing perturbations to enhance hallucinations, fails to effectively enlarge the logits gap between hallucinatory and targeted tokens, resulting in unsatisfactory outputs. On the contrary, our method addresses the issue by significantly amplifying the logits gap between hallucinatory and targeted tokens. (Right) The performance of various methods on CHAIR metrics. Our HIO generates descriptions with fewer hallucination tokens compared to other visual contrastive decoding methods, achieving lower scores on the CHAIRs and CHAIRi metrics.
@@
However, the inherent uncontrollable nature of global visual uncertainty challenges the precise induction of hallucinatory tokens. This limitation significantly undermines the effectiveness of these methods in reducing hallucinations and may inadvertently lead to undesired hallucinatory outputs. As shown in the left portion of the Fig. 1 ‘Spoon’, ‘Table’, and ‘Fork’ are identified as hallucinated words, while ’People’ being the accurate term. For Greedy Decoding method shown in Fig. 1 (a), ‘Table’ is selected as the final output based on the logits distribution. Moreover, although Visual Contrastive Decoding introduces perturbations to images to enhance hallucinations in Fig. 1 (v), it fails to widen the logits gaps between hallucinatory and targeted tokens, yielding a new hallucination as ‘Fork’.
To tackle this issue, we conducted the theoretical analysis to explore mechanisms for more effective contrast decoding (refer to Section 5 for detailed information on the process). Theoretically, a clear distinction between hallucinatory and target tokens can significantly enhance the effectiveness of contrast decoding methods in mitigating hallucinations. Based on this crucial insight, we introduce a novel optimization strategy called Hallucination-Induced Optimization (HIO). This strategy enhances the distinction between hallucinatory and targeted tokens by utilizing a refined theoretical preference model(as shown in the Fig. 1 on the left, section (c)), accurately outputting the correct result, ‘People’. Consequently, this improves the efficiency of contrast decoding, thereby mitigating hallucinations in Large Vision-Language Models (LVLMs). Furthermore, our proposed method significantly reduces hallucinations in LVLMs compared to existing contrast decoding methods(as shown in the Fig. 1 on the right). To sum up, our main contributions are as follows:
@@
    We conducted a comprehensive theoretical analysis to explore mechanisms that enhance the effectiveness of the contrast decoding process.
@@ 2 Related Work
Before the advent of Large Language Models (LLMs), "hallucination" in natural language processing (NLP) primarily referred to generating nonsensical or source-deviating content Lee et al. (2018); Zhou et al. (2020); Lin et al. (2021); Ji et al. (2023); Zhang et al. (2023); Shi et al. (2023). Recent studies have tackled the complexities of object hallucination in Large Vision-Language Models (LVLMs), focusing on evaluation and detection methods Wang et al. (2023a); Liu et al. (2023a); Li et al. (2023b); Lovenia et al. (2023). The CHAIR metric Rohrbach et al. (2018) evaluates the exact match between generated and ground-truth image captions, while POPE Li et al. (2023b) assesses the model’s awareness of object existence through binary classification.
@@
The decoding method determines the generation of text tokens at each time step within language models. Traditional decoding strategies such as beam search Boulanger-Lewandowski et al. (2013), top-k decoding Fan et al. (2018), and sampling methods Holtzman et al. (2019), despite their widespread use, are prone to producing hallucinatory content. Recent research Li et al. (2022); Chuang et al. (2023); Leng et al. (2023); Huang et al. (2023) has made attempts to address this issue by proposing better decoding methods. For instance, Leng et al. (2023) uses contrastive decoding in LVLMs; However, global visual uncertainty poses challenges to the precise induction of hallucinatory tokens, limiting the effectiveness of mitigation strategies and risking unwanted hallucinations. To address this, we developed Hallucination-Induced Optimization (HIO), a novel strategy that enhances the contrast between hallucinatory and targeted tokens. Fig.1 presents the comparison results, where our approach demonstrates superior performance than other decoding methods.
@@ 3 Preliminaries
We first review the Contrast Decoding pipeline in Leng et al. (2023) (and later Zhang et al. (2024); Favero et al. (2024)). Then take a close look at the Bradley-Terry model Bradley and Terry (1952) and its application such as Direct Preference Optimization Rafailov et al. (2024). Inspired by these studies, we propose our Hallucination-Induced Optimization.
@@
where $\mathcal{D}$ is a dataset of prompts and $\alpha$ is a coefficient to control KL-divergence term. However, optimizing the above loss term with common strategies like proximal policy optimization (PPO) Schulman et al. (2017) is complex to tune. Recently, direct preference optimization (DPO) Rafailov et al. (2024) simplifies the above process by leveraging preference data for optimization. Here, the preference data is defined as $\mathcal{D}=\{x^{(i)},y_{w}^{(i)},y_{l}^{(i)}\}_{i=1}^{N}$, where $y_{w}^{(i)}$ and $y_{l}^{(i)}$ represent preferred and dispreferred responses given an input prompt $x$. These are then presented to human labelers who express preferences for one answer, denoted as $y_{w}\succ y_{l}\mid x$ where $y_{w}$ and $y_{l}$ denotes the preferred and dispreferred respectively. Following a Bradley-Terry model Bradley and Terry (1952), the probably of obtaining each preference pair is:
@@ 4 Method
An overview of the proposed HIO method is shown in Fig. 3. It constructs a more-hallucinated LVLM by inducing hallucinations from the original LVLM to amplify the contrast between hallucinatory and targeted tokens, thereby enhancing the efficiency of contrast decoding and mitigating hallucinations in LVLMs. In Section 4.1, we harness a fine-tuned theoretical preference model to amplify the contrast between hallucinatory and targeted tokens. Furthermore, to induce more potential hallucinations for effective contrast decoding, we propose to amplify multiple hallucination tokens based on a theoretical foundation in Section 4.2. Section 4.3 introduces additional constraints to overcome the limitations of existing classification loss in amplifying the contrast between hallucinatory and targeted tokens.
@@ 4.1 Contrary Bradley-Terry Model (CBTM)
We harness a fine-tuned theoretical preference model (i.e., Contrary Bradley-Terry Model Bradley and Terry (1952)) to amplify the contrast between hallucinatory and targeted tokens. The studies on hallucination mitigation Zhao et al. (2023); Yu et al. (2023); Zhou et al. (2024) utilize BT model by defining the non-hallucinatory output as $y_{w}$ and the hallucinatory output as $y_{l}$. Subsequently, they employ BT model training to incentivize the model to prioritize outputs without hallucinations over those containing them.
However, within the context of contrast decoding, inducing hallucinations is crucial, and the resulting model output must satisfy the criteria outlined in Eqn. (17). (The detailed derivation of this formula is provided in Section5). To meet the requirements specified in Eqn. (17), the logits associated with hallucinated tokens ${\hat{l_{i}}^{\{v,x,y_{<t}\}}}$ need amplification, while at least one of the logits for the correct token ${\hat{l_{j}}^{\{v,x,y_{<t}\}}}$ must be reduced. In contrast to the prevailing research efforts focused on alleviating hallucinations, our approach enables the model to learn to fit the distribution containing hallucinations while avoiding convergence with the distribution of correct outputs. The details are outlined as follows. To regulate ${\hat{l_{i}}^{\{v,x,y_{<t}\}}}$ and ${\hat{l_{j}}^{\{v,x,y_{<t}\}}}$, we utilize the dataset introduced by Yu et al. (2023). This dataset is notable for providing a pair of outputs per input, with the output paragraphs being mostly identical except for differences in certain words or short phrases. By leveraging this dataset, we approximate the conditions outlined in Eqn. (17) within a unified statement. Different from Eqn. (5), we apply the Bradley-Terry (BT) Bradley and Terry (1952) model in a reversed way, the objective is:
@@ 4.2 Amplification of Multiple Targeted Hallucination (AMTH)
Multiple Hallucination-Induced Optimization. Achieving the desired distribution through single positive and negative sample fitting preference training is not feasible, leading conventional Direct Preference Optimization (DPO) applications Zhao et al. (2023); Yu et al. (2023); Zhou et al. (2024) to overlook a significant number of hallucinations. Thus, drawing inspiration from the implications of Eqn. (17), our approach strategically induces multiple hallucinations to increase the probability of producing a correct word in the output. As demonstrated in Eqn. (17), effective contrast decoding necessitates not only the amplification of one hallucination but also the consideration of a diverse set of potential hallucinations. We propose the simultaneous fitting of multiple pairs of preference data when modeling distributions for the same input preference, treating all pairs of preference data with equal importance. Based on Eqn. (6), we apply the Bradley-Terry (BT) Bradley and Terry (1952) model in a multi-pair way, the objective is:
@@ 4.3 Advanced Constraints for Inducing (ACI)
where $m$ represents the length of the sentence $y_{l}$ and $\{k_{T}\}$ is token of each word, and the definition of $\hat{l_{i}}^{\{v,x,y_{<t}\}}$ is shown in Section 5. While the use of cross-entropy to minimize encoding length helps the model align with the desired output sentence, it does not consistently ensure that the logits of amplified hallucinations meet the conditions specified in Eqn. (17).
For example, the goal of Eqn. (8) is to increase $\pi_{\theta}(y_{l}|v,x)$, but both increasing $\exp{\hat{l_{k_{t}}}^{\{v,x,y_{<t}\}}}$ or decreasing $\sum_{j}^{N}\exp{\hat{l_{j}}^{\{v,x,y_{<t}\}}}$ can achieve this goal. Meanwhile, decreasing the value of $\sum_{j}^{N}\exp{\hat{l_{j}}^{\{v,x,y_{<t}\}}}$ can also allow $\pi_{\theta}(y_{w}|v,x)$ to meet the optimization criteria. As shown in Fig. 3, the blue discount, representing the disparity between the logits of the hallucination token and the correct token, typically exhibits a positive trend. Nevertheless, it’s important to note occasional segments where this value dips below zero. To tackle this issue, we further add restrictions based on Eqn. (8):
@@
5 Fundamental Requirements for Contrast Decoding
@@
where $\delta^{{}^{\prime}\{v,x,y_{<t}\}}$ denotes the result of the subtraction between the logits of all hallucinated vocabulary tokens and the logits after their ideal amplification. $\delta^{{}^{\prime}\{v,x,y_{<t}\}}$ represents the outcome of the subtraction between the logits corresponding to all correct vocabulary tokens and the logits under the ideal scenario. Upon expansion of the left side of the equation, the following result is obtained:
@@
where $J$ represents $\frac{(1+\alpha)}{\alpha}\sum_{i=k_{1}}^{k_{m}}({l_{i}^{\{v,x,y_{<t}\}}}-l_{j}^{\{v,x,y_{<t}\}})$. In the context of the contrast decoding method, given that the parameters of the original model remain invariant, the output can be characterized as a constant. Equation  17 delineates the logits for all hallucinated tokens ${\hat{l_{i}}^{\{v,x,y_{<t}\}}}$ and contrasts these with the logits of a single correct token ${\hat{l_{j}}^{\{v,x,y_{<t}\}}}$. It postulates that, for an optimal logits output, a pronounced divergence must be maintained between the logits of hallucinated tokens and the logit of the correct token.

### v2

@@ Abstract
Although Large Visual Language Models (LVLMs) have demonstrated exceptional abilities in understanding multimodal data, they invariably suffer from hallucinations, leading to a disconnection between the generated text and the corresponding images. Almost all current visual contrastive decoding methods attempt to mitigate these hallucinations by introducing visual uncertainty information that appropriately widens the contrastive logits gap between hallucinatory and targeted ones. However, due to uncontrollable nature of the global visual uncertainty, they struggle to precisely induce the hallucinatory tokens, which severely limits their effectiveness in mitigating hallucinations and may even lead to the generation of undesired hallucinations. To tackle this issue, we conducted the theoretical analysis to promote the effectiveness of contrast decoding. Building on this insight, we introduce a novel optimization strategy named Hallucination-Induced Optimization (HIO). This strategy seeks to amplify the contrast between hallucinatory and targeted tokens relying on a fine-tuned theoretical preference model (i.e., Contrary Bradley-Terry Model), thereby facilitating efficient contrast decoding to alleviate hallucinations in LVLMs. Extensive experimental research demonstrates that our HIO strategy can effectively reduce hallucinations in LVLMs, outperforming state-of-the-art methods across various benchmarks. Code is released at https://github.com/BT-C/HIO.
@@ 1 Introduction
The recent success of Large Vision-Language Models (LVLMs) marks a major milestone in artificial intelligence research (OpenAI, 2023; Alayrac et al., 2022; Li et al., 2023b; Liu et al., 2023e; Zhu et al., 2023b; Bai et al., 2023b; Dai et al., 2023; Wang et al., 2023c; Driess et al., 2023). By seamlessly integrating visual cues with Large Language Models (LLMs), LVLMs have demonstrated unparalleled expertise in multimodal comprehension, logical reasoning, and interactive engagement. This integration has ushered in a new era in AI, breaking through traditional limitations and enabling a more holistic understanding of complex information OpenAI (2023); Yang et al. (2023); Lu et al. (2023); Yuan et al. (2022); Sun et al. (2024). Despite these advancements, certain challenges remain, particularly the issue of hallucination Li et al. (2023d); Gunjal et al. (2023); Liu et al. (2023b); Lovenia et al. (2023). Hallucination occurs when the language model generates content that deviates from the image’s actual content, including imagined objects, fabricated scenes, incorrect spatial relationships, and misidentified categories.
Figure 1: (Left) Challenges and Solutions of Contrast Decoding Strategy. Visual Contrastive Decoding, despite introducing perturbations to induce hallucinations, fails to effectively enlarge the logits gap between hallucinatory and targeted tokens, resulting in unsatisfactory outputs. On the contrary, our method addresses the issue by significantly amplifying the logits gap between hallucinatory and targeted tokens. (Right) The performance of various methods on CHAIR metrics. Our HIO generates descriptions with fewer hallucination tokens compared to other visual contrastive decoding methods, achieving lower scores on the CHAIRs and CHAIRi metrics.
@@
However, the inherent uncontrollable nature of global visual uncertainty challenges the precise induction of hallucinatory tokens. This limitation significantly undermines the effectiveness of these methods in reducing hallucinations and may inadvertently lead to undesired hallucinatory outputs. As shown in the left portion of the Fig. 1 Spoon, Table, and Fork are identified as hallucinated words, while People being the accurate term. For Greedy Decoding method shown in Fig. 1 (a), Table is selected as the final output based on the logits distribution. Moreover, although Visual Contrastive Decoding introduces perturbations to images to enhance hallucinations in Fig. 1 (b), it fails to widen the logits gaps between hallucinatory (Spoon, Table, and Fork) and targeted tokens (People), yielding a new hallucination as Fork.
To tackle this issue, we conducted the theoretical analysis to explore mechanisms for more effective contrast decoding (refer to Section 5 for detailed information on the process). Theoretically, a clear distinction between hallucinatory and target tokens can significantly enhance the effectiveness of contrast decoding methods in mitigating hallucinations. Based on this crucial insight, we introduce a novel optimization strategy called Hallucination-Induced Optimization (HIO). This strategy enhances the distinction between hallucinatory and targeted tokens by utilizing a refined theoretical preference model(as shown in the Fig. 1 on the left, section (c)), accurately outputting the correct result, People. Consequently, this improves the efficiency of contrast decoding, thereby mitigating hallucinations in Large Vision-Language Models (LVLMs). Furthermore, our proposed method significantly reduces hallucinations in LVLMs compared to existing contrast decoding methods(as shown in the Fig. 1 on the right). To sum up, our main contributions are as follows:
@@
    We conducted a comprehensive theoretical analysis to explore mechanisms that enhance the effectiveness of the contrast decoding strategy.
@@ 2 Related Work
Before the advent of Large Language Models (LLMs), "hallucination" in natural language processing (NLP) primarily referred to generating nonsensical or source-deviating content Lee et al. (2018); Zhou et al. (2020); Lin et al. (2021); Ji et al. (2023); Zhang et al. (2023b); Shi et al. (2023b). Recent studies have tackled the complexities of object hallucination in Large Vision-Language Models (LVLMs), focusing on evaluation and detection methods Wang et al. (2023a); Liu et al. (2023a); Li et al. (2023d); Lovenia et al. (2023). The CHAIR metric Rohrbach et al. (2018) evaluates the exact match between generated and ground-truth image captions, while POPE Li et al. (2023d) assesses the model’s awareness of object existence through binary classification.
@@
The decoding method determines the generation of text tokens at each time step within language models. Traditional decoding strategies such as beam search Boulanger-Lewandowski et al. (2013), top-k decoding Fan et al. (2018), and sampling methods Holtzman et al. (2019), despite their widespread use, are prone to producing hallucinatory content. Recent research Li et al. (2022b); Chuang et al. (2023); Leng et al. (2023); Huang et al. (2023) has made attempts to address this issue by proposing better decoding methods. For instance, Leng et al. (2023) uses contrastive decoding in LVLMs; However, global visual uncertainty poses challenges to the precise induction of hallucinatory tokens, limiting the effectiveness of mitigation strategies and risking unwanted hallucinations. To address this, we developed Hallucination-Induced Optimization (HIO), a novel strategy that enhances the contrast between hallucinatory and targeted tokens. Fig.1 presents the comparison results, where our approach demonstrates superior performance than other decoding methods.
@@ 3 Preliminaries
We first review the Contrast Decoding pipeline in Leng et al. (2023) (and later Zhang et al. (2024); Favero et al. (2024)). Then take a close look at the Bradley-Terry model Bradley, Terry (1952) and its application such as Direct Preference Optimization Rafailov et al. (2024). Inspired by these studies, we propose our Hallucination-Induced Optimization.
@@
where $\mathcal{D}$ is a dataset of prompts and $\alpha$ is a coefficient to control KL-divergence term. However, optimizing the above loss term with common strategies like proximal policy optimization (PPO) Schulman et al. (2017) is complex to tune. Recently, direct preference optimization (DPO) Rafailov et al. (2024) simplifies the above process by leveraging preference data for optimization. Here, the preference data is defined as $\mathcal{D}=\{x^{(i)},y_{w}^{(i)},y_{l}^{(i)}\}_{i=1}^{N}$, where $y_{w}^{(i)}$ and $y_{l}^{(i)}$ represent preferred and dispreferred responses given an input prompt $x$. These are then presented to human labelers who express preferences for one answer, denoted as $y_{w}\succ y_{l}\mid x$ where $y_{w}$ and $y_{l}$ denote the preferred and dispreferred respectively. Following a Bradley-Terry model (Bradley, Terry, 1952), the probability of obtaining each preference pair is:
@@ 4 Method
An overview of the proposed HIO method is shown in Fig. 2. It constructs a more-hallucinated LVLM by inducing hallucinations from the original LVLM to amplify the contrast between hallucinatory and targeted tokens, thereby enhancing the efficiency of contrast decoding and mitigating hallucinations in LVLMs. In Section 4.1, we harness a fine-tuned theoretical preference model to amplify the contrast between hallucinatory and targeted tokens. Furthermore, to induce more potential hallucinations for effective contrast decoding, we propose to amplify multiple hallucination tokens based on a theoretical foundation presented in Eqn. 17 of Section 5. This theory demonstrates that effective contrastive decoding requires a consistent difference between the logits of potential hallucinated tokens and the correct token. And Section 4.3 introduces additional constraints to overcome the limitations of existing classification loss in amplifying the contrast between hallucinatory and targeted tokens.
@@ 4.1 Contrary Bradley-Terry Model (CBTM)
We harness a fine-tuned theoretical preference model (i.e., Contrary Bradley-Terry Model (Bradley, Terry, 1952)) to amplify the contrast between hallucinatory and targeted tokens. The studies on hallucination mitigation Zhao et al. (2023); Yu et al. (2023); Zhou et al. (2024) utilize BT model by defining the non-hallucinatory output as $y_{w}$ and the hallucinatory output as $y_{l}$. Subsequently, they employ BT model training to incentivize the model to prioritize outputs without hallucinations over those containing them.
However, within the context of contrast decoding, inducing hallucinations is crucial, and the resulting model output must satisfy the criteria outlined in Eqn. (17). (The detailed derivation of this formula is provided in Section5). To meet the requirements specified in Eqn. (17), the logits associated with hallucinated tokens ${\hat{l_{i}}^{\{v,x,y_{<t}\}}}$ need amplification, while at least one of the logits for the correct token ${\hat{l_{j}}^{\{v,x,y_{<t}\}}}$ must be reduced. In contrast to the prevailing research efforts focused on alleviating hallucinations, our approach enables the model to learn to fit the distribution containing hallucinations while avoiding convergence with the distribution of correct outputs. The details are outlined as follows. To regulate ${\hat{l_{i}}^{\{v,x,y_{<t}\}}}$ and ${\hat{l_{j}}^{\{v,x,y_{<t}\}}}$, we utilize the dataset introduced by Yu et al. (2023). This dataset is notable for providing a pair of outputs per input, with the output paragraphs being mostly identical except for differences in certain words or short phrases. By leveraging this dataset, we approximate the conditions outlined in Eqn. (17) within a unified statement. Different from Eqn. (5), we apply the Bradley-Terry (BT) (Bradley, Terry, 1952) model in a reversed way, the objective is:
@@ 4.2 Amplification of Multiple Targeted Hallucination (AMTH)
Multiple Hallucination-Induced Optimization. Achieving the desired distribution through single positive and negative sample fitting preference training is not feasible, leading conventional Direct Preference Optimization (DPO) applications Zhao et al. (2023); Yu et al. (2023); Zhou et al. (2024) to overlook a significant number of hallucinations. Thus, drawing inspiration from the implications of Eqn. (17), our approach strategically induces multiple hallucinations to increase the probability of producing a correct word in the output. As demonstrated in Eqn. (17), effective contrast decoding necessitates not only the amplification of one hallucination but also the consideration of a diverse set of potential hallucinations. We propose the simultaneous fitting of multiple pairs of preference data when modeling distributions for the same input preference, treating all pairs of preference data with equal importance. Based on Eqn. (6), we apply the Bradley-Terry (BT) (Bradley, Terry, 1952) model in a multi-pair way, the objective is:
@@ 4.3 Advanced Constraints for Inducing (ACI)
where $m$ represents the length of the sentence $y_{l}$ and $\{k_{T}\}$ is token of each word, and the definition of $\hat{l_{i}}^{\{v,x,y_{<t}\}}$ is shown in Section 5. While the use of cross-entropy to minimize encoding length helps the model align with the desired output sentence, it does not consistently ensure that the logits of induced hallucinations meet the conditions specified in Eqn. (17).
For example, the goal of Eqn. (8) is to increase $\pi_{\theta}(y_{l}|v,x)$, but both increasing $\exp{\hat{l_{k_{t}}}^{\{v,x,y_{<t}\}}}$ or decreasing $\sum_{j}^{N}\exp{\hat{l_{j}}^{\{v,x,y_{<t}\}}}$ can achieve this goal. Meanwhile, decreasing the value of $\sum_{j}^{N}\exp{\hat{l_{j}}^{\{v,x,y_{<t}\}}}$ can also allow $\pi_{\theta}(y_{w}|v,x)$ to meet the optimization criteria. As shown in Fig. 3, the blue curve, representing the disparity between the logits of the hallucinatory and targeted tokens, typically exhibits a positive trend. Nevertheless, it’s important to note occasional segments where this value dips below zero. To tackle this issue, we further add restrictions based on Eqn. (8):
@@
5 Fundamental Conditions for Contrast Decoding
@@
where $\delta^{{}^{\prime}\{v,x,y_{<t}\}}$ denotes the result of the subtraction between the logits of all hallucinated vocabulary tokens and the logits after their ideal amplification. $\delta^{*\{v,x,y_{<t}\}}$ represents the outcome of the subtraction between the logits corresponding to all correct vocabulary tokens and the logits under the ideal scenario. Eqn. 14 represents a theoretical upper bound, which guides us in enhancing the effectiveness of Contrast Decoding method for hallucination elimination by ensuring that the logits of all hallucinated words are lower than those of the correct words. Upon expansion of the left side of the equation, the following result is obtained:
@@
where $J$ represents $\frac{(1+\alpha)}{\alpha}\sum_{i=k_{1}}^{k_{m}}({l_{i}^{\{v,x,y_{<t}\}}}-l_{j}^{\{v,x,y_{<t}\}})$. In the context of the contrast decoding method, given that the parameters of the original model remain invariant, the output can be characterized as a constant. Eqn. 17 delineates the logits for all hallucinated tokens ${\hat{l_{i}}^{\{v,x,y_{<t}\}}}$ and contrasts these with the logits of a single correct token ${\hat{l_{j}}^{\{v,x,y_{<t}\}}}$. It postulates that, for an optimal logits output, a pronounced divergence must be maintained between the logits of hallucinated tokens and the logit of the correct token.
Eqn. 17 illustrates that hallucinations can be effectively eliminated through contrastive decoding if the difference between the logits of the hallucinatory token and the correct token in the ‘Evil’ LVLM’s output (Left part of Eqn.17) exceeds that in the original LVLM output ($J$ in Eqn.17). For example, as depicted in the lower part of Fig. 2, where "Dogs" is a hallucination and "Benches" is the correct label, the hallucination of "Dogs" is removed when the difference between the logits for "Dogs" and "Benches" in the ‘Evil’ LVLM output surpasses the difference in the original LVLM output. When this condition is met for all potential hallucinations, all hallucinations are effectively eliminated.